OpenStudy (anonymous):

3. The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line? (1 point)

OpenStudy (anonymous):

(-5,0),(0,-2) m=(y2-y1)/(x2-x) solve for m 1st

OpenStudy (anonymous):

wanna try it?

OpenStudy (anonymous):

I got -0.4

OpenStudy (anonymous):

correct y - y1 = m(x - x1) plug in the numbers u have m= -0.4 (x1,y1)=(-5,0)

OpenStudy (anonymous):

Wait how do I do this?

OpenStudy (anonymous):

what is the value for y1? what is the value for x1? what is the value for m?

OpenStudy (anonymous):

y1 is the y coordinate of one of your points (you choose) x1 is the x coordingat of the SAME point m is the gradient of the line

OpenStudy (anonymous):

coordinate*

OpenStudy (anonymous):



OpenStudy (anonymous):

Theresthe answer choices

OpenStudy (anonymous):

y - 0 = -0.4*(x - -5) simplify this